OpenStudy (anonymous):

i got a ACT test study and i need help with some of it reply if u are willing to help

OpenStudy (anonymous):

car A averages 16 miles a gallon car B average 35 miles a gallon at this rate how much more gallons will car A and B need to make a 1120 mile trip

OpenStudy (phi):

if you multiply 1120 miles by the ratio of gallons/mile , you will get gallons in other words, car A will use \( 1120\ miles \cdot \frac{1\ gallon}{16 miles} \) and car B will use \( 1120\ miles \cdot \frac{1\ gallon}{35 \ miles} \) and the difference is \[ 1120 \cdot \frac{1}{16 }- 1120 \cdot \frac{1}{35 } \]

OpenStudy (anonymous):

thx so helpful

OpenStudy (phi):

if you don't have a calculator you can do the division in steps (it helps to know some "rules" see https://en.wikipedia.org/wiki/Divisibility_rule) for example, 16 is 4*4 divide 4 into 1120 to get 280 then divide that by 4 to get 70 we have 70 - 1120/(5*7)

OpenStudy (anonymous):

the intergers below adds up to 390 x-3 x-2 x-1 x x+1 what is the value of x

OpenStudy (phi):

if you add them up (use algebra) what do you get ?

OpenStudy (anonymous):

-55

OpenStudy (phi):

first add up the x's. how many x's do you have?

OpenStudy (anonymous):

5

OpenStudy (phi):

for example x+x is 2x when you add up the x's you get 5x (not just 5) now add up the numbers. what do you get ?

OpenStudy (anonymous):

5x-3-2-1+1?

OpenStudy (phi):

yes, but you can simplify the numbers and get just 1 total for the numbers

OpenStudy (phi):

in other words do -3-2 -1+1

OpenStudy (anonymous):

-5?

OpenStudy (phi):

yes. all of that means the sum of the "x stuff" is 5x-5 and they say they add up to 390 we can say 5x-5= 390 add 5 to both sides 5x-5+5 = 390+5 and simplify. what do you get ?

OpenStudy (phi):

on the left side you have -5+5 which adds to 0 so 5x+0 on the left but adding 0 does nothing, so we can just say 5x= 390+5 what do you get on the right side?

OpenStudy (anonymous):

x=79?

OpenStudy (phi):

yes

OpenStudy (phi):

5x= 395 divide both sides by 5 x= 395/5 = 79

OpenStudy (anonymous):

ight thanks and is it ok if i message u tomarrow for more help if i need it

OpenStudy (phi):

yes